You had to choose 4 dogs at random. There are 6 dogs that are breed A, 8 dogs that are breed B, and 12 that are breed C. Given that the 2 breed C were chosen, what are the chances of picking at least 1 breed A.
Answer #1

Given that 2 breed C were chosen, the probability of picking at least 1 breed A is computed here as:

= 1 - Probability that there is no A in the remaining 2 selected

= 1 - (Number of ways to select remaining 2 from 8 breed B dogs / Total ways to select 2 from the remaining 14 dogs of breed A or B)

= 1 - (8c2 / 14c2)

= 1 - (28 / 91)

= 0.6923

Therefore 0.6923 is the required probability here.
Answer #2

SOLUTION :


If 2 dogs chosen are C in the first 2 pickings , then we are left with A = 6, B = 8 and C = 10 , Total dogs left = 24 

P(at least 1 dog is A from next two pickings I  2 dogs are C)

=  1 - P( no dog A in 3rd and 4th picking | 2 dogs are C)

= 1 - 18/24 * 17/23

= 0.44 57 (ANSWER).

answered by: Tulsiram Garg